OpenStudy (anonymous):

An object is dropped from 33 feet below the pinnacle atop a 1477-ft building. The height (h) of the object after t seconds is given by the equation h=-16t^2+1444. How many seconds until the object is on the ground?

OpenStudy (anonymous):

@phuong.tong

OpenStudy (anonymous):

@Squirrels @Wolfgirl12 @quickstudent @vzfreakz @jazzyspazzy

OpenStudy (anonymous):

@Roxeyrai73 @sleepyjess @TheEdwardsFamily @brittanydosey @Mindblast3r @Donblue22 @PlanSlam @punksauce

OpenStudy (anonymous):

@hecnev @HotPinkGirl @winsteria @EclipsedStar @elizabethroland53 @Ellie202000

OpenStudy (anonymous):

@Abmon98 @aum @abb0t @Jasmine456 @jdoe0001

OpenStudy (kendricklamar2014):

The height is 0 when it hits the ground so we do: h=-16t^2+1444 = 0

OpenStudy (kendricklamar2014):

Which equals: 16^2 = 1,444

OpenStudy (anonymous):

ok, so we have to figure out what t is in order for h to be 0?

OpenStudy (kendricklamar2014):

Now we do: t^2 = 1,444/16

OpenStudy (anonymous):

is that the next step?

OpenStudy (anonymous):

@KendrickLamar2014 are you there??

OpenStudy (kendricklamar2014):

Positive Solution: T = 38/4 = 19/2 = 9.5 seconds

OpenStudy (kendricklamar2014):

Thats how you solve this problem @Lucinda12345

OpenStudy (kendricklamar2014):

Yes

OpenStudy (anonymous):

ok, sorry my computer is super slow

OpenStudy (anonymous):

Thanks for the help

OpenStudy (kendricklamar2014):

No Problem